# Toponium: the smallest bound state and simplest hadron in quantum mechanics

### v1

@@ Abstract
We explore toponium, the smallest known quantum bound state of a top quark and its antiparticle, bound by the strong force. With a Bohr radius of $8\times 10^{-18}$ m and a lifetime of $2.5\times 10^{-25}$ s, toponium uniquely probes microphysics. Unlike all other hadrons, it is governed by ultraviolet freedom, exhibiting feeble interactions at distances much smaller than $10^{-15}$ m, rather than infrared slavery that characterizes powerful interactions at approximately $10^{-15}$ m. This distinction offers novel insights into quantum chromodynamics. Our analysis reveals a toponium signal exceeding five standard deviations in the distribution of the cross section ratio between $e^{+}e^{-}\rightarrow b\bar{b}$ and $e^{+}e^{-}\rightarrow q\bar{q}$ ($q=b,~{}c,~{}s,~{}d,~{}u$), based on 400 fb-1 (1 fb = $10^{-43}~{}{\rm m}^{2}$) of data collected at around 341 GeV, driven by quantum interference. This discovery enables a top quark mass measurement with an uncertainty reduced by a factor of ten compared to current precision levels. Detection prospects at the Circular Electron Positron Collider or the Future Circular Lepton Collider underscore their potential to revolutionize our understanding of quantum mechanics.
Introduction
@@
The top quark ($t$), the heaviest known element particle, plays a pivotal role in Standard Model (SM) of particle physics. Its unique properties, particularly its mass, have far-reaching implications for our understanding of fundamental forces and the stability of the universe. Toponium, a bound state of a top quark and an anti-top quark ($\bar{t}$), represents the smallest quantum mechanical bound state known to date. Formed by strong force, toponium has nearly twice the mass of a single top quark [1, 2], making it about 360 times heavier than a proton and twice as massive as a gold nucleus. The Bohr radius of toponium, $8\times 10^{-18}$ m, is significantly smaller than those of hydrogen ($5.3\times 10^{-11}$ m) and proton (0.83 fm, 1 fm=$10^{-15}$ m). With an extremely short lifetime of $2.5\times 10^{-25}$ s, the study of toponium offers a unique opportunity to explore fundamental aspects of quantum mechanics, including entanglement [3, 4, 5], as well as to explore quantum gravity and extra dimensions [6, 7] at scales of $10^{-18}$ m and $10^{-25}$ s.
@@
The top quark mass ($m_{t}$) is a crucial fundamental parameter in SM, as $m_{t}$ significantly influences our understanding of SM, particularly its vacuum stability, and provides a window into potential new physics beyond SM [8, 9, 10, 11]. Despite its importance, $m_{t}$ cannot be predicted theoretically and must be measured experimentally. The most precise measurement to date [2] yields an on-shell mass of $m_{t}^{\text{OS}}=(172.52\pm 0.14_{\rm stat.}\pm 0.30_{\rm syst.})~{}{\rm GeV}/c^{2}=(172.52\pm 0.33)~{}{\rm GeV}/c^{2}$ (GeV, gigaelectronvolts; $c$, the speed of light in vacuum, is ignored in subsequent units for simplicity). While the statistical uncertainty (${\rm stat.}$) can be reduced with longer data acquisition time, the systematic uncertainty (${\rm syst.}$) defines the ultimate precision of the measurement.
@@
The most precise method for determining $m_{t}$ involves the near-threshold production of $t\bar{t}$ at future electron positron ($e^{+}e^{-}$) colliders, such as the Circular Electron Positron Collider (CEPC) [12, 13] and the Future Circular Lepton Collider (FCC-ee) [14]. Projections suggest that $m_{t}$ could be measured to a precision of $25-59$ MeV at CEPC [15] and $40-75$ MeV at FCC-ee [16], with statistical uncertainties as low as 9 MeV at integrated luminosities of $100-200$ fb-1 (1 fb = $10^{-43}~{}{\rm m}^{2}$) at around 343 GeV, although systematic uncertainties will dominate. These uncertainties arise from multiple sources, such as beam energy calibration, theoretical modeling, and detector performance [17, 15]. Therefore, the accurate determination of $m_{t}$ relies heavily on the development of novel method and advanced systematic control.
@@
Furthermore, toponium distinguishes itself among hadrons because its behavior is governed by ultraviolet freedom, rather than infrared slavery of quantum chromodynamics (QCD) at distances of roughly $10^{-15}$ m [27, 28]. This unique characteristic enables toponium to be theoretically described by perturbative QCD (PQCD), thereby offering new insights into confinement within the Yang-Mills field. Given its high binding energy, the phase transition between toponium and free top quark pairs could have profound implications for the structure of the quark-gluon plasma [29], cooling rates [30], and the supercooled electroweak phase transition in the early universe [31, 32, 33].
In this study, we propose a method to discover $J_{t}$ and accurately determine $m_{t}^{\text{1S}}$ by measuring the cross section ratio:
@@
at $e^{+}e^{-}$ colliders near $2m_{t}^{\text{1S}}$, where $b$ is the bottom quark and $q$ is a quark for $u$$d$$s$$c$, or $b$. Future measurements of $R_{b}$ at the Z pole are expected to achieve precision levels of $4\times 10^{-5}$ at CEPC [12] and less than $6\times 10^{-5}$ at FCC-ee [34]. The relative precision of $R_{b}$ at $365$ GeV with $1.5~{}{\rm ab}^{-1}$ data at FCC-ee will be $9.6\times 10^{-4}$ [35], which results an absolute precision level of $1.4\times 10^{-4}$. The other possible methods to identify toponium can be found in Refs. [36, 37] (Further details can be found in Supplemental material.).
Modeling $J_{t}$ Interaction
We introduce the Lagrangian density $\mathcal{L}_{J_{t}}$ of the field $J_{t}$ into that of SM, $\mathcal{L}_{\text{SM}}$ [36] as follows (Supplemental material):
@@
where the coupling factors $g^{V}_{{J_{t}}ff}$ and $g^{A}_{{J_{t}}ff}$ are determined by non-relativistic QCD [38], $f$ represents the field of either the $b$ quark or the electron, $\gamma_{\mu}$ and $\gamma^{5}$ are the Dirac gamma matrices, and ${J_{t}}^{\mu}$ is the field of $J_{t}$.
The ${J_{t}}$ bound state is obtained by the Schrödinger equation with the Coulomb potential between $t$ and $\bar{t}$ as a lattice QCD result at an infinite quark mass [39, 40] (Supplemental material),
$V(r)=-\frac{\hbar c\lambda}{r},$ (3)
where $\hbar$ is the reduced Planck constant, the Coulombic coefficient $\lambda=0.285\pm 0.011$ [39, 40], and $r$ is the distance between $t$ and $\bar{t}$. We derive the binding energy $B_{J_{t}}$, mass $m_{J_{t}}=2m_{t}^{\text{OS}}-B_{J_{t}}=2m_{t}^{\text{1S}}$, and the square of the wave function at the origin $|\psi_{J_{t}}(0)|^{2}$:
@@
The $B_{J_{t}}$ and $|\psi_{J_{t}}(0)|^{2}$ have also been calculated in PQCD, but the convergence of the perturbative series up to next-to-next-to-next-to leading order (NNNLO) is still unsatisfactory (Supplemental material), requiring the consideration of higher order corrections. Including the Higgs and photon contributions, the NNNLO binding energy is 3.194 GeV, which is comparable to the $B_{J_{t}}$ in Eq.(Modeling $J_{t}$ Interaction) (Supplemental material). The estimated production cross section of spin-singlet toponium at the Large Hadron Collider (LHC) at 13 TeV, assuming identical masses and wave functions for the toponium states, is 6.24 pb (Supplemental material). This estimate agrees with the value of 6.43 pb reported in Refs. [36, 37] and the recent CMS result of $(7.1\pm 0.8)$ pb [41].
The width $\Gamma_{J_{t}}$ of $J_{t}$ is dominated by twice the width of the top quark [42] (Supplemental material):
$\Gamma_{J_{t}}=(2.603\pm 0.019)~{}{\rm GeV}.$ (5)
@@
where $m_{J_{t}}$ and $\Gamma_{J_{t}}$ are its mass and width, $\lambda^{3}$ is part of the coupling factors $g^{V}_{{J_{t}}ff}$ and $g^{A}_{{J_{t}}ff}$, and $\sqrt{s}$ is center of mass energy. The real part of the denominator $s-m_{J_{t}}^{2}$ goes from negative to 0 to positive as $\sqrt{s}$ goes over $m_{J_{t}}$. In the amplitudes ${\mathcal{M}}(\gamma)$ and ${\mathcal{M}}(Z)$ corresponding to the contributions of the photon (with $m_{\gamma}=\Gamma_{\gamma}=0$) and the Z boson (with $m_{Z}=91.188~{}$ GeV) in SM respectively, the sign of the term $s-m^{2}$ remains unaltered. This results in a characteristic dip-bump structure in the $R_{b}$ distribution, stemming from the fact that the production probabilities for quantum interference are proportional to the square of the sum of amplitudes, namely $|{\mathcal{M}}(J_{t})+{\mathcal{M}}(\gamma)+{\mathcal{M}}(Z)|^{2}$. This dip-bump structure in the $R_{b}$ distribution of SM with $J_{t}$ is seen in Fig. 1, while the $R_{b}$ distribution of SM without $J_{t}$ is very close to a horizontal straight line. By observing the deviation of $R_{b}$ distribution from the straight line, we can identify $J_{t}$, and a precise measurement of this deviation allows us to determine $m_{t}^{\text{1S}}$.
@@
Measuring $R_{b}$ and identifying the signature of $J_{t}$
The $R_{b}$, an observable which can be precisely measured in a pseudo-experiment with the CEPC scenario, is determined by [44] (Supplemental material):
$R_{b}$ $=$ $\frac{f_{\rm ISR}^{udsc}\left(N_{\text{tag}}\varepsilon_{\text{hadron}}^{udsc}-N_{\text{hadron}}\varepsilon_{\text{tag}}^{{udsc}}\right)}{N_{\text{tag}}\left(f_{\rm ISR}^{udsc}\varepsilon_{\text{hadron}}^{udsc}-f_{\rm ISR}^{b}\varepsilon^{b}_{\text{hadron}}\right)+N_{\text{hadron}}\left(f_{\rm ISR}^{b}\varepsilon_{\text{tag}}^{b}-f_{\rm ISR}^{udsc}\varepsilon_{\text{tag}}^{udsc}\right)}.$ (7)
Here, the $\varepsilon^{q}_{\text{hadron}}$ ($\varepsilon^{q}_{\text{tag}}$) is the detection efficiency (double $b$ tagged efficiency) for $q=u,~{}d,~{}s,~{}c,~{}b$ after the event selection criteria mentioned in supplementary materials, and the $\varepsilon_{\text{hadron}}^{udsc}$ ($\varepsilon_{\text{tag}}^{udsc}$) is the averaged detection efficiency (double $b$ tagged efficiency) weighted by the cross section of $e^{+}e^{-}\to q\bar{q}~{}(q=u,~{}d,~{}s,~{}c)$, where $\varepsilon_{\text{hadron}}^{{udsc}}={52.48\%}$, $\varepsilon_{\text{hadron}}^{{b}}={52.70\%}$, $\varepsilon_{\text{tag}}^{b}={46.42\%}$, and $\varepsilon_{\text{tag}}^{{udsc}}={0.04\%}$ (Supplemental material). Meanwhile, the $N_{\text{tag}}$ $(N_{\rm hadron})$ is the number of selected double $b$ tagged hadronic events (the number of hadronic events) from the experiment datasets, and can be obtained with information of the cross section, integrated luminosity, and detection efficiency at an energy point (Supplemental material). The $f^{b}_{\rm ISR}$ and $f^{udsc}_{\rm ISR}$ are the correction factors that include the initial state radiation (ISR) and energy spread (Supplementary Material). Then the $R_{b}$ value and its corresponding statistical uncertainty can be determined at an energy point.
@@
$\chi^{2}_{\rm Model}=\sum_{i}{\left(\frac{R_{b}^{\text{PEX}}(\mathcal{L}_{i},\sqrt{s_{i}})-R_{b}^{\rm Model}(\mathcal{L}_{i},\sqrt{s_{i}};m_{J_{t}},\lambda)}{\Delta R_{b}^{\text{PEX}}(\mathcal{L}_{i},\sqrt{s_{i}})}\right)^{2}}.$ (8)
The $\mathcal{L}_{i}$ is the integrated luminosity at $\sqrt{s_{i}}$, and ${\Delta R_{b}^{\text{PEX}}(\mathcal{L}_{i},\sqrt{s_{i}})}$ is the uncertainty of $R_{b}^{\text{PEX}}(\mathcal{L}_{i},\sqrt{s_{i}})$. The fitting function $R_{b}^{\rm Model}(\mathcal{L}_{i},\sqrt{s_{i}};m_{J_{t}},\lambda)$ gives a theoretical $R_{b}$ value at an energy point, where free parameters $m_{J_{t}}$ and $\lambda$ can be varied to minimize $\chi^{2}_{\rm Model}$.
@@
Figure 1: The blue dashed line represents $R_{b}$ distribution in SM without $J_{t}$ model, while the red solid line represents $R_{b}$ distribution in SM with $J_{t}$. Additionally, the statistical significance of three energy points is indicated. The insert shows the LO Feynman diagram for the process $e^{+}e^{-}\rightarrow q\bar{q}$, with the corresponding propagator represented by a photon, a Z boson, or a $J_{t}$.
@@
The main sources of systematic uncertainties in the $R_{b}$ measurements are the energy scale of $\sqrt{s}$, energy spread, detection efficiency, and backgrounds. All systematic uncertainties are more than an order of magnitude smaller than the statistical uncertainties (Supplementary Material), leading to a $J_{t}$ signal significance of $6.25\sigma$ even when considering the total systematic uncertainties.
When the true values of $m_{J_{t}}$ and $\lambda$ vary within their respective $1\sigma$ ranges, we determine the corresponding significance of the $J_{t}$ signal. This significance consistently exceeds the $5\sigma$ threshold and is shown in Fig. 2. A significance level greater than $5\sigma$ indicates that the probability of SM with $J_{t}$ is incorrect compared to SM without $J_{t}$ being correct is less than $5.7\times 10^{-7}$.
Extracting the top quark mass $m_{t}^{\text{1S}}$
The top quark mass $m_{t}^{\text{1S}}$ (along with the Coulombic coefficient $\lambda$) is extracted by minimizing $\chi^{2}_{\text{SM with}~{}J_{t}}$. The statistical uncertainty $({\Delta m_{t}^{\text{1S}}})_{\rm stat.}$ is found to be $\rm 32.2~{}MeV$, primarily determined by the second energy point $\sqrt{s_{2}}$. The systematic uncertainties, $({\Delta m_{t}^{\text{1S}}})_{\rm syst.}$, arise from several sources, including the energy scale and energy spread, detection efficiency, backgrounds, SM parameters, higher order corrections, and others (luminosity measurements and backgrounds from two-photon and $WWZ$ processes).
@@
Experimentally, the background, as well as the detection efficiency and the weak mixing angle $\theta_{\rm W}$, can be determined primarily by the data from $Z$, $WW$, and $HZ$ peaks directly translated to the $m_{J_{t}}$ region [46]. The detection efficiency related uncertainty in $R_{b}$ is ${\varepsilon_{\text{hadron}}^{udsc}}/{\varepsilon_{\text{tag}}^{b}}$ (Supplemental material), which is consistent that in the $R_{b}$ measurement at the Z pole. After including uncertainties from Monte Carlo (MC) statistics and energy-momentum requirement, the uncertainty associated with the detection efficiency in $R_{b}$ is $19.1\times 10^{-6}$ (Supplemental material). This is about half of the total uncertainty of the $R_{b}$ measurement at the Z-pole at CEPC, which is $4\times 10^{-5}$ [12]. Consequently, this contributes 1.3 MeV to $\Delta m_{t}^{\text{1S}}$.
By comparing the background contributions estimated in theory and observed in experiment, a relative precision of $(\alpha_{s}/\pi)^{2}\sim 10^{-3}$ can be achieved, contributing 0.9 MeV to $\Delta m_{t}^{\text{1S}}$. Additionally, $\sin^{2}\theta_{\rm W}$ can attain an accuracy of $1\times 10^{-5}$ [12], thereby contributing 0.4 MeV to $\Delta m_{t}^{\text{1S}}$. The NNLO correction of $R_{b}$ yields a contribution of 0.1 MeV to $\Delta m_{t}^{\text{1S}}$. The contribution of other sources, such as luminosity and backgrounds from two-photon and $WWZ$ processes, is less than 0.2 MeV for $\Delta m_{t}^{\text{1S}}$ (Supplemental material).
@@
Table 1: Systematic uncertainties of the $m_{t}$ measurement in MeV. The “Other” category includes luminosity measurements and backgrounds from two-photon and $WWZ$ processes.
@@
SM parameters ($\theta_{\rm W},~{}\alpha_{s}$) 3.2 17 $0.4$
@@
The upcoming $e^{+}e^{-}$ colliders CEPC and FCC-ee aim to achieve integrated luminosities of 1000 fb-1 [13] and 1700 fb-1 [14], respectively, near the $t\bar{t}$ production threshold. With these luminosities, the systematic uncertainty on $m_{t}^{\text{1S}}$ is expected to remain at 1.9 MeV for both CEPC and FCC-ee. If the luminosities are chosen as $\mathcal{L}=(80,~{}900,~{}20)~{}{\rm fb}^{-1}$ and $\mathcal{L}=(80,~{}1600,~{}20)~{}{\rm fb}^{-1}$ for CEPC and FCC-ee at (340.076, 341.598, 342.771) GeV, respectively, our results show that, the statistical uncertainty, $({\Delta m_{t}^{\text{1S}}})_{\rm stat.}$, is estimated to be 18.6 MeV for CEPC and 13.9 MeV for FCC-ee, or 11.2 MeV when data from both colliders are combined. A detailed comparison of the top quark mass measurement methods at CEPC [15], FCC-ee [16], and this work is presented in Table 2.
Table 2: Detailed comparison of the top quark mass measurement methods: final states, signal and background cross sections, and uncertainty estimation at CEPC [15], FCC-ee [16], and this work.

### v2

@@ Abstract
We explore toponium, the smallest known quantum bound state of a top quark and its antiparticle, bound by the strong force. With a Bohr radius of $8\times 10^{-18}$ m and a lifetime of $2.5\times 10^{-25}$ s, toponium uniquely probes microphysics. Unlike all other hadrons, it is governed by ultraviolet freedom. This distinction offers novel insights into quantum chromodynamics. Our analysis reveals a toponium signal exceeding $5\sigma$ in the distribution of the cross section ratio between $e^{+}e^{-}\rightarrow b\bar{b}$ and $e^{+}e^{-}\rightarrow q\bar{q}$ ($q=b,c,s,d,u$), based on 400 fb-1 of data collected at $\sqrt{s}\approx 341~{}{\rm GeV}$. This discovery enables a top quark mass measurement with an uncertainty reduced by a factor of ten compared to current precision levels. Moreover, this method improves the systematic uncertainty by at least a factor of 12 compared to any other possible methods.
@@
Introduction—The top quark ($t$), the heaviest known element particle, plays a pivotal role in Standard Model (SM) of particle physics. Its unique properties, particularly its mass, have far-reaching implications for our understanding of fundamental forces and the stability of the universe. Toponium, a bound state of a top quark and an anti-top quark ($\bar{t}$), represents the smallest quantum mechanical bound state known to date. Formed by strong force, toponium has nearly twice the mass of a single top quark [1, 2]. With the smallest Bohr radius of $8\times 10^{-18}$ m and an extremely short lifetime of $2.5\times 10^{-25}$ s, the study of toponium offers a unique opportunity to explore fundamental aspects of quantum mechanics, including entanglement [3, 4, 5, 6], as well as to explore quantum gravity and extra dimensions [7, 8] at scales of $10^{-18}$ m and $10^{-25}$ s.
@@
The top quark mass ($m_{t}$) is a crucial fundamental parameter in SM, as $m_{t}$ significantly influences our understanding of SM, particularly its vacuum stability, and provides a window into potential new physics beyond SM [9, 10, 11, 12]. Despite its importance, $m_{t}$ cannot be predicted theoretically and must be measured experimentally. The most precise measurement to date [2] yields an on-shell mass of $m_{t}^{\text{OS}}=(172.52\pm 0.14_{\rm stat.}\pm 0.30_{\rm syst.})~{}{\rm GeV}=(172.52\pm 0.33)~{}{\rm GeV}$. While the statistical uncertainty (${\rm stat.}$) can be reduced with longer data acquisition time, the systematic uncertainty (${\rm syst.}$) defines the ultimate precision of the measurement.
@@
The most precise method for determining $m_{t}$ involves the near-threshold production of $t\bar{t}$ at future electron positron ($e^{+}e^{-}$) colliders, such as the Circular Electron Positron Collider (CEPC) [13, 14] and the Future Circular Lepton Collider (FCC-ee) [15]. Projections suggest that $m_{t}$ could be measured to a precision of $25-59$ MeV at CEPC [16] and $40-75$ MeV at FCC-ee [17], with statistical uncertainties as low as 9 MeV at integrated luminosities of $100-200$ fb-1 around 343 GeV, although systematic uncertainties will dominate. These uncertainties arise from multiple sources, such as beam energy calibration, theoretical modeling, and detector performance [18, 16]. Therefore, the accurate determination of $m_{t}$ relies heavily on the development of novel method and advanced systematic control.
@@
Furthermore, toponium distinguishes itself among hadrons because its behavior is governed by ultraviolet freedom, rather than infrared slavery of quantum chromodynamics (QCD) at distances of roughly $200~{}{\rm MeV}$ [28, 29]. This unique characteristic enables toponium to be theoretically described by perturbative QCD (PQCD), thereby offering new insights into confinement within QCD [28]. Given its high binding energy, the phase transition between toponium and free top quark pairs could have profound implications for the structure of the quark-gluon plasma [30], cooling rates [31], and the supercooled electroweak phase transition in the early universe [32, 33, 34].
In this Letter, we propose a method to discover $J_{t}$ and accurately determine $m_{t}^{\text{1S}}$ by measuring the cross section ratio:
@@
at $e^{+}e^{-}$ colliders near $2m_{t}^{\text{1S}}$, where $b$ is the bottom quark and $q$ is a quark for $u$$d$$s$$c$, or $b$. Future measurements of $R_{b}$ at the $Z$ pole are expected to achieve precision levels of $4\times 10^{-5}$ at CEPC [13] and less than $6\times 10^{-5}$ at FCC-ee [35]. The relative precision of $R_{b}$ at $365$ GeV with $1.5~{}{\rm ab}^{-1}$ data will be $9.6\times 10^{-4}$ at FCC-ee [36], which results an absolute precision level of $1.4\times 10^{-4}$. The other possible methods to identify toponium can be found in Refs. [37, 38, 39] and related papers.
Modeling $J_{t}$ Interaction—We introduce the Lagrangian density $\mathcal{L}_{J_{t}}$ of the field $J_{t}$ into that of SM, $\mathcal{L}_{\text{SM}}$ [37] as follows:
@@
where the coupling factors $g^{V}_{{J_{t}}ff}$ and $g^{A}_{{J_{t}}ff}$ are determined by non-relativistic QCD (NRQCD) [40], $f$ represents the field of either the $b$ quark or the electron, and ${J_{t}}^{\mu}$ is the field of $J_{t}$.
The ${J_{t}}$ bound state is obtained by the Schrödinger equation with the Coulomb potential between $t$ and $\bar{t}$ as a lattice QCD result at an infinite quark mass [41, 42],
$V(r)=-\frac{\lambda}{r}.$ (3)
The Coulombic coefficient $\lambda=0.285\pm 0.011$ [41, 42], and $r$ is the distance between $t$ and $\bar{t}$. We derive the binding energy $B_{J_{t}}$, mass $m_{J_{t}}=2m_{t}^{\text{OS}}-B_{J_{t}}=2m_{t}^{\text{1S}}$, and the square of the wave function at the origin $|\psi_{J_{t}}(0)|^{2}$:
@@
The $B_{J_{t}}$ and $|\psi_{J_{t}}(0)|^{2}$ have also been calculated in PQCD, but the convergence of the perturbative series up to next-to-next-to-next-to-leading order (NNNLO) is still unsatisfactory (End Matter), requiring the consideration of higher order corrections. Including the Higgs and photon contributions [43], the NNNLO binding energy is 3.194 GeV, which is comparable to the $B_{J_{t}}$ in Eq.(Toponium: the smallest bound state and simplest hadron in quantum mechanics). The estimated production cross section of spin-singlet toponium at the Large Hadron Collider (LHC) at 13 TeV, assuming identical masses and wave functions for the toponium states, is 6.24 pb. This estimate agrees with the value of 6.43 pb reported in Refs. [37, 38] and the recent CMS result of $(7.1\pm 0.8)$ pb [44].
The width $\Gamma_{J_{t}}$ of $J_{t}$ is dominated by twice the width of the top quark $\Gamma_{t}$ [45, 46]:
$\Gamma_{J_{t}}$ $=$ $2\Gamma_{t}\left(1-\frac{\lambda^{2}}{8}\right)+\Gamma_{\rm Anni}=(2.603\pm 0.019)~{}{\rm GeV},$ (5)
where the coefficient $\left(1-\lambda^{2}/8\right)$ comes from corrections for the fact that bound top quarks are different from free top quarks [47]. The annihilation decay width of $J_{t}$, $\Gamma_{\rm Anni}=0.0093~{}{\rm GeV}$, is calculated using NRQCD [40].
@@
where $m_{J_{t}}$ and $\Gamma_{J_{t}}$ are its mass and width, $\lambda^{3}$ is part of the coupling factors $g^{V}_{{J_{t}}ff}$ and $g^{A}_{{J_{t}}ff}$, and $\sqrt{s}$ is the center of mass energy. The real part of the denominator $s-m_{J_{t}}^{2}$ goes from negative to 0 to positive as $\sqrt{s}$ goes over $m_{J_{t}}$. In the amplitudes ${\mathcal{M}}(\gamma)$ and ${\mathcal{M}}(Z)$ corresponding to the contributions of the photon (with $m_{\gamma}=\Gamma_{\gamma}=0$) and the $Z$ boson (with $m_{Z}=91.188~{}$ GeV) in SM respectively, the sign of the term $s-m^{2}$ remains unaltered. This results in a characteristic dip-bump structure in the $R_{b}$ distribution, stemming from the fact that the production probabilities for quantum interference are proportional to the square of the sum of amplitudes, namely $|{\mathcal{M}}(J_{t})+{\mathcal{M}}(\gamma)+{\mathcal{M}}(Z)|^{2}$. This dip-bump structure in the $R_{b}$ distribution of SM with $J_{t}$ is seen in Fig. 1, while the $R_{b}$ distribution of SM without $J_{t}$ is very close to a horizontal straight line. By observing the deviation of $R_{b}$ distribution from the straight line, we can identify $J_{t}$, and a precise measurement of this deviation allows us to determine $m_{t}^{\text{1S}}$.
@@
Measuring $R_{b}$—The $R_{b}$, an observable which can be precisely measured in a pseudo-experiment with the CEPC scenario, is obtained by [49] (End Matter):
$R_{b}=\frac{N_{\text{tag}}F_{\text{hadron}}^{udsc}-N_{\text{hadron}}F_{\text{tag}}^{{udsc}}}{N_{\text{tag}}(F_{\text{hadron}}^{udsc}-F^{b}_{\text{hadron}})+N_{\text{hadron}}(F_{\text{tag}}^{b}-F_{\text{tag}}^{udsc})}.$ (7)
Here, $F=\varepsilon f_{\rm ISR}$, the detection efficiency for $b$ quarks after event selection criteria is denoted by $\varepsilon^{b}_{\text{hadron}}$ (for hadronic events) and $\varepsilon^{b}_{\text{tag}}$ (for double $b$-tagged events). Similarly, the averaged detection efficiency for light quarks ($q=u,d,s,c$) is denoted by $\varepsilon_{\text{hadron}}^{udsc}$ (for hadronic events) and $\varepsilon_{\text{tag}}^{udsc}$ (for double $b$-tagged events), weighted by the cross section of $e^{+}e^{-}\to q\bar{q}$. The values are $\varepsilon_{\text{hadron}}^{udsc}=52.48\%$, $\varepsilon_{\text{hadron}}^{b}=52.70\%$, $\varepsilon_{\text{tag}}^{b}=46.42\%$, and $\varepsilon_{\text{tag}}^{udsc}=0.04\%$ (End Matter). The correction factors $f^{udsc}_{\rm ISR}$ and $f^{b}_{\rm ISR}$ are used to model the effects of initial state radiation (ISR) and energy spread. Meanwhile, $N_{\text{tag}}$ $(N_{\rm hadron})$ is the number of selected double $b$-tagged hadronic events (the number of hadronic events) from the experiment datasets, and can be obtained with information of the cross section, integrated luminosity, and detection efficiency at an energy point. If we ignore the small values $\varepsilon_{\text{hadron}}^{udsc}-\varepsilon_{\text{hadron}}^{b}$, $\varepsilon_{\text{tag}}^{udsc}$, and $1-{f_{\rm ISR}^{b}}/{f_{\rm ISR}^{udsc}}$, we can get $R_{b}\sim(N_{\text{tag}}\varepsilon_{\text{hadron}}^{udsc})/(N_{\text{hadron}}\varepsilon_{\text{tag}}^{b})$. Then the $R_{b}$ value and its corresponding statistical uncertainty can be determined at an energy point (End Matter).
@@
$\chi^{2}_{\rm Model}=\sum_{i}{\left(\frac{R_{b}^{\text{PEX}}(\mathcal{L}_{i},\sqrt{s_{i}})-R_{b}^{\rm Model}(\mathcal{L}_{i},\sqrt{s_{i}})}{\Delta R_{b}^{\text{PEX}}(\mathcal{L}_{i},\sqrt{s_{i}})}\right)^{2}}.$ (8)
The $\mathcal{L}_{i}$ is the integrated luminosity at $\sqrt{s_{i}}$, and ${\Delta R_{b}^{\text{PEX}}(\mathcal{L}_{i},\sqrt{s_{i}})}$ is the uncertainty of $R_{b}^{\text{PEX}}(\mathcal{L}_{i},\sqrt{s_{i}})$. The fitting function $R_{b}^{\rm Model}(\mathcal{L}_{i},\sqrt{s_{i}})$ gives a theoretical $R_{b}$ value at an energy point, where free parameters $m_{J_{t}}$ and $\lambda$ can be varied to minimize $\chi^{2}_{\rm Model}$.
@@
Figure 1: The blue dashed line represents $R_{b}$ distribution in SM without $J_{t}$, the red solid line represents $R_{b}$ distribution in SM with $J_{t}$, and the statistical significance of the three energy points is indicated. The upper left insert shows $R_{b}$ values for $m_{t}^{\rm Input}$ and $m_{t}^{\rm Input}\pm 30~{}\text{MeV}$ near the second energy point. The lower right insert shows the LO Feynman diagram for $e^{+}e^{-}\rightarrow q\bar{q}$, where the propagator can be a photon, a $Z$, or a $J_{t}$.
@@
The main sources of systematic uncertainties in the $R_{b}$ measurements are the energy scale of $\sqrt{s}$, energy spread, detection efficiency, and backgrounds. All systematic uncertainties are more than an order of magnitude smaller than the statistical uncertainties (End Matter), leading to a $J_{t}$ signal significance of $6.25\sigma$ even when the total systematic uncertainties are taken into account.
When the true values of $m_{J_{t}}$ and $\lambda$ vary within their respective $1\sigma$ ranges, we determine the corresponding significance of the $J_{t}$ signal. This significance consistently exceeds the $5\sigma$ threshold and is shown in Fig. 2.
Extracting $m_{t}^{\text{1S}}$—The top quark mass $m_{t}^{\text{1S}}$ (along with the coefficient $\lambda$) is extracted by minimizing $\chi^{2}_{\text{SM with}~{}J_{t}}$. The statistical uncertainty $({\Delta m_{t}^{\text{1S}}})_{\rm stat.}$ is found to be $\rm 32.2~{}MeV$, primarily determined by the second energy point $\sqrt{s_{2}}$. The systematic uncertainties, $({\Delta m_{t}^{\text{1S}}})_{\rm syst.}$, arise from several sources, including the energy scale and energy spread, detection efficiency, SM parameters, higher order corrections, and others (luminosity measurements and backgrounds from two-photon and $WWZ$ processes).
@@
Experimentally, the backgrounds, as well as the detection efficiency and the weak mixing angle $\theta_{\rm W}$, can be determined primarily by the data from $Z$, $WW$, and $HZ$ peaks, which are directly translated to the $m_{J_{t}}$ region [51]. The detection efficiency related uncertainty in $R_{b}$ is ${\varepsilon_{\text{hadron}}^{udsc}}/{\varepsilon_{\text{tag}}^{b}}$, which is consistent that in the $R_{b}$ measurement at the $Z$ pole. After including uncertainties from Monte Carlo statistics and energy-momentum requirement, the uncertainty associated with the detection efficiency in $R_{b}$ is $19.1\times 10^{-6}$ (End Matter). This is about half of the total uncertainty of the $R_{b}$ measurement at the $Z$ pole at CEPC, which is $4\times 10^{-5}$ [13]. Consequently, this contributes 1.3 MeV to $\Delta m_{t}^{\text{1S}}$.
By comparing the background contributions estimated in theory and observed in experiment, a relative precision of $(\alpha_{s}/\pi)^{2}\sim 10^{-3}$ can be achieved, contributing 0.9 MeV to $\Delta m_{t}^{\text{1S}}$. Additionally, $\sin^{2}\theta_{\rm W}$ can attain an accuracy of $1\times 10^{-5}$ [13], thereby contributing 0.4 MeV to $\Delta m_{t}^{\text{1S}}$. The NNLO correction of $R_{b}$ yields a contribution of 0.1 MeV to $\Delta m_{t}^{\text{1S}}$. The contribution of other sources, such as luminosity and backgrounds from two-photon and $WWZ$ processes, is less than 0.2 MeV for $\Delta m_{t}^{\text{1S}}$ (End Matter).
@@
Table 1: Systematic uncertainties of the $m_{t}$ measurement in MeV. The “Other" category includes luminosity measurements and backgrounds from two-photon and $WWZ$ processes.
@@
SM parameters ($\theta_{\rm W},\alpha_{s}$) 3.2 17 0.4
@@
The upcoming $e^{+}e^{-}$ colliders aim to achieve integrated luminosities of 1,000 fb-1 at CEPC [14] and 1,700 fb-1 at FCC-ee [15], respectively, near the $t\bar{t}$ production threshold. With these luminosities, the systematic uncertainty on $m_{t}^{\text{1S}}$ is expected to remain at 1.9 MeV for both CEPC and FCC-ee. If the luminosities are chosen as $\mathcal{L}=(80,900,20)~{}{\rm fb}^{-1}$ and $\mathcal{L}=(80,1,\!600,20)~{}{\rm fb}^{-1}$ for CEPC and FCC-ee at (340.076, 341.598, 342.771) GeV, respectively, our results show that, the statistical uncertainty, $({\Delta m_{t}^{\text{1S}}})_{\rm stat.}$, is estimated to be 18.6 MeV for CEPC and 13.9 MeV for FCC-ee, or 11.2 MeV when data from both colliders are combined. A detailed comparison of the top quark mass measurement methods at CEPC [16], FCC-ee [17], and in this work is presented in Table 2.
Table 2: Detailed comparison of the top quark mass measurement methods: final states, signal and background cross sections, and uncertainty estimation at CEPC [16], FCC-ee [17], and in this work.